# Unlearning via Sparse Representations

### v1

@@ E.1 Training Details and Hyperparameters for training the original DKVB Models
CIFAR-10 CIFAR-100 LACUNA-100
top-k 1 10 10
Key Dimension 8 8 8
# of Key Init Epochs 10 10 10
Type of Value Init Gaussian Random Zeros Uniform Random
# of Codebooks 256 256 256
# of Key-Value Pairs per Codebook 4096 4096 4096
Optimizer Adam Adam Adam
LR 0.1 0.3 0.3
Batch Size 256 256 256
Epochs 74 71 7
E.2 Training Details and Hyperparameters for training the original Baseline Models
For the baseline models, we deliberately train them to similar test ($\mathcal{D}_{test}$) accuracies as the models with a Discrete Key Value Bottleneck to ensure a fair comparison for unlearning. Table 5 shows the hyperparameters used for training the baseline models
@@ E.2 Training Details and Hyperparameters for training the original Baseline Models
CIFAR-10 CIFAR-100 LACUNA-100
LR 0.001 0.01 0.01
Batch Size 256 256 256
Epochs 2 2 2
E.3 Training Details and Hyperparameters for SCRUB
For the baseline, we run SCRUB on the model with linear layer. One epoch consists of one min step and may or may not contain a max step. One max step is included in every epoch for the first msteps epochs. We tune the hyperparameter msteps in our experiments and pick the case where the model is able to best recover its performance on the retain set test data and consider this model as the final unlearned model.We mention the hyperparameters used for running SCRUB for corresponding to the results presented in Section 5.3 in Table 6. In this case, training of SCRUB is stopped when either (a) the forget set accuracy has dropped to 0% without damaging the retain set accuracy or (b) at the end of 10 epochs. In the latter case, the value of the retain set test accuracy reported is the highest value in the training when under the constraint that the forget set test accuracy is 0 at that point. Results presented in Section 5.4 also uses the same set of hyperparameters except min-step which is always 10 since we train all the methods for 10 epochs. We do not consider cases where SCRUB is not able to completely unlearn the forget class within the given computational budget.
Table 6: Hyperparameters for SCRUB + Linear Layer Experiments shown in Section 5.3
CIFAR-10 CIFAR-100 LACUNA-100
Forget Set Batch Size 256 256 256
Retain Set Batch Size 256 256 256
# of max-steps (msteps) 3 9 5
# of min-steps 3 10 7
LR 0.001 0.01 0.01
Optimizer Adam Adam Adam
Batch Size 256 256 256
E.4 Training Details and Hyperparameters for Retraining Experiments
Once the DKVB models are unlearned using Unlearning via Activations and Unlearning via Examples, we retraining them in order to make a fair comparison with the baseline. Thus, during retraining, the initial performance of these models on the retain set is same as the final performance of the zero-shot unlearned models. Table 7 show the hyperparameters used for retraining the unlearned DKVB models.
@@ E.4 Training Details and Hyperparameters for Retraining Experiments
Appendix F Comparison of Runtimes
In this section, we present a comparison of runtimes of our approach against the baseline, i.e. SCRUB (Kurmanji et al., 2023) as a proxy for comparing the compute requirements of the two approaches. Table 8 compares the runtimes of the proposed approaches against the baseline. We observe that the runtime for the proposed approaches is orders smaller than that of the baseline.
Table 8: Comparison of runtimes between the proposed methods and the baseline on CIFAR-10, CIFAR-100 and LACUNA-100.
CIFAR-10 CIFAR-100 LACUNA-100
Runtimes (in seconds)
DKVB via Activations (sec 5.2) 5.02 1.57 2.74
DKVB via Examples (sec 5.2) 13.98 6.65 13.03
Linear Layer + SCRUB 288.80 921.56 553.31
Appendix G Experiments with a ResNet Backbone
To demonstrate the the proposed approaches are agnostic to the choice of the backbone, we run the same set of experiments presented in Section 5.2 and Section 5.3 on CIFAR-10 with an ImageNet supervised pretrained backbone. Figure 7 shows the scatter plot of retain set test performance vs. forget set test performance on CIFAR-10 with a ResNet backbone. Table 9 shows the comparison against baseline.
Table 9: Comparison between the proposed methods with a pretrained ResNet50 backbone and the baseline on CIFAR-10. We compare the change in performance on the retain and forget test data relative to the originally trained models. For the baseline, we report two cases: Case B where the model unlearns the forget set completely but the retain set performance is not preserved very well, and Case A where the retain set performance is not preserved very well, but the model does not unlearn the forget set completely
CIFAR-10
Rel. change from base model $\mathcal{D}_{test}^{retain}$ $\mathcal{D}_{test}^{forget}$
DKVB via Activations (sec 5.2) 0.04% -100%
DKVB via Examples (sec 5.2) -0.07% -100%
Linear Layer + SCRUB (A) 0.44% -96.04%
Linear Layer + SCRUB (B) -11.75% -100%
(a) CIFAR-10 (via Activations)
(b) CIFAR-10 (via Examples)
Figure 7: Performance on the retain set test data vs. Performance on the forget set test data for DKVB models with a supervised pretrained ResNet50 backbone on CIFAR-10.
There is a steep but drop in the retain set test accuracy towards the end when $N_{a}$ and $N_{e}$ are high enough. However, speaking in absolute terms, the drop seems to be not significant.
Appendix H Experimental Results on ImageNet-1k
Table 10: Comparison between the proposed methods and the baselines on ImageNet-1k. We compare the change in performance on the retain and forget test data relative to the originally trained models.
Rel. change from base model $\mathcal{D}_{test}^{retain}$ $\mathcal{D}_{test}^{forget}$
DKVB via Activations (sec 5.2) 0.15% -100%
DKVB via Examples (sec 5.2) -0.03% -100%
Linear Layer + SCRUB 7.31% -100%
(a) ImageNet (via Activations)
(b) ImageNet (via Examples)
Figure 8: Performance on the retain set test data vs. Performance on the forget set test data for DKVB models on ImageNet-1k
pPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@ A.7.1 Training Details and Hyperparameters for training the original DKVB Models
Backbone Hyperparameter CIFAR-10 CIFAR-100 LACUNA-100 ImageNet-1k
ViT/B-32 top-k 1 10 10 1
Key Dimension 8 8 8 14
# of Key Init Epochs 10 10 10 10
Type of Value Init Gaussian Random Zeros Uniform Random Zeros
# of Codebooks 256 256 256 256
# of Key-Value Pairs per Codebook 4096 4096 4096 4096
Optimizer Adam Adam Adam Adam
LR 0.1 0.3 0.3 0.3
Batch Size 256 256 256 256
Epochs 74 71 7 3
ResNet-50 top-k 1 2 1 1
Key Dimension 14 14 8 14
# of Key Init Epochs 10 10 10 10
Type of Value Init Zeros Random Gaussian Random Gaussian Random
# of Codebooks 256 256 256 256
# of Key-Value Pairs per Codebook 4096 4096 4096 4096
Optimizer Adam Adam Adam Adam
LR 0.3 0.3 0.1 0.3
Batch Size 256 256 256 256
Epochs 70 4 1 5
A.7.2 Training Details and Hyperparameters for training the original Baseline Models
For the baseline models, we deliberately train them to similar test ($\mathcal{D}_{test}$) accuracies as the models with a Discrete Key Value Bottleneck to ensure a fair comparison for unlearning. Table 8 shows the hyperparameters used for training the baseline models.
@@ A.7.2 Training Details and Hyperparameters for training the original Baseline Models
Backbone Hyperparameter CIFAR-10 CIFAR-100 LACUNA-100 ImageNet-1k
ViT/B-32 LR 0.001 0.01 0.01 0.01
Batch Size 256 256 256 512
Epochs 1 7 13 1
ResNet-50 LR 0.01 0.001 0.01 0.001
Batch Size 256 256 512 512
Epochs 2 72 73 11
A.7.3 Training Details and Hyperparameters for SCRUB
For the baseline, we run SCRUB on the model with linear layer. One epoch consists of one min step and may or may not contain a max step. Hence the values of min steps and epochs are always same. One max step is included in every epoch for the first msteps epochs. We tune the hyperparameter msteps in our experiments and pick the case where the model is able to best recover its performance on the retain set test data and consider this model as the final unlearned model. We mention the hyperparameters used for running SCRUB corresponding to the results presented in Section 5.2.1 in Table 9. In this case, training of SCRUB is stopped when the forget set accuracy has either dropped to 0% or converged at a close to 0% value without damaging the retain set accuracy. Results presented in Appendix A.5 also use the same set of hyperparameters except min-step which is always 10 since we train all the methods for 10 epochs.
Table 9: Hyperparameters for SCRUB + Linear Layer Experiments shown in Section 5.2.1
Backbone Hyperparameter CIFAR-10 CIFAR-100 LACUNA-100 ImageNet-1k
ViT/B-32 Forget Set Batch Size 256 256 256 512
Retain Set Batch Size 256 256 256 512
# of max-steps (msteps) 3 9 5 3
# of min-steps / # of epochs 3 10 7 3
LR 0.001 0.01 0.01 0.001
Optimizer Adam Adam Adam Adam
ResNet-50 Forget Set Batch Size 256 256 256 512
Retain Set Batch Size 256 256 256 512
# of max-steps (msteps) 9 3 3 3
# of min-steps / # of epochs 10 30 30 10
LR 0.01 0.001 0.01 0.001
Optimizer Adam Adam Adam Adam
A.7.4 Training Details and Hyperparameters for Retraining Experiments
Once the DKVB models are unlearned using Unlearning via Activations and Unlearning via Examples, we retraining them in order to make a fair comparison with the baseline. Thus, during retraining, the initial performance of these models on the retain set is same as the final performance of the unlearned models. Table 10 show the hyperparameters used for retraining the unlearned DKVB models.